# Fast or Better? Balancing Accuracy and Cost in Retrieval-Augmented Generation with Flexible User Control

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a powerful approach to mitigate large language model (LLM) hallucinations by incorporating external knowledge retrieval. However, existing RAG frameworks often apply retrieval indiscriminately,leading to inefficiencies—over-retrieving when unnecessary or failing to retrieve iteratively when required for complex reasoning. Recent adaptive retrieval strategies, though adaptively navigates these retrieval strategies, predict only based on query complexity and lacks user-driven flexibility, making them infeasible for diverse user application needs. In this paper, we introduce a novel user-controllable RAG framework that enables dynamic adjustment of the accuracy-cost trade-off. Our approach leverages two classifiers: one trained to prioritize accuracy and another to prioritize retrieval efficiency. Via an interpretable control parameter $\alpha$, users can seamlessly navigate between minimal-cost retrieval and high-accuracy retrieval based on their specific requirements. We empirically demonstrate that our approach effectively balances accuracy, retrieval cost, and user controllability  [^1] Code can be found at https://github.com/JinyanSu1/Flare-Aug., making it a practical and adaptable solution for real-world applications.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Khandelwal et al. (2019); Izacard et al. (2023) has emerged as a promising approach to address large language models (LLM) hallucinations, outputs that appear accurate but are actually factually incorrect. Hallucinations more often occur when the facts necessary to answer a query are outside the scope of the model’s training either because the facts are too recent, too obscure or answered by proprietary data Trivedi et al. (2022a); Yao et al. (2022). By integrating retrieval modules, RAG enables LLMs to access and incorporate relevant external information, helping them stay updated with evolving world knowledge and reduce hallucination.
@@
Early work on retrieval-augmented LLMs primarily focused on single-hop queries Lazaridou et al. (2022); Ram et al. (2023), where a single retrieval step retrieves relevant information based solely on the query. However, many complex queries require multi-step reasoning to arrive at the correct answer. For instance, a question such as, ”When did the people who captured Malakoff come to the region where Philipsburg is located?”, cannot be answered with a single retrieval step. Instead, such queries necessitate an iterative retrieval process, where the model retrieves partial knowledge, reasons over it, and conducts additional retrievals based on intermediate conclusions Trivedi et al. (2022a); Press et al. (2022); Jeong et al. (2024). While multi-step retrieval enables deeper reasoning, it also incurs significant computational overhead. Conversely, for simple queries, LLMs may already encode sufficient knowledge to generate an accurate response without retrieval. However, many existing RAG frameworks apply retrieval indiscriminately, either over-relying on external knowledge when parametric memory suffices or failing to retrieve iteratively when deeper reasoning is required. This lack of adaptability results in inefficiencies: unnecessary retrieval increases latency and computational cost, while insufficient retrieval leads to incomplete or incorrect answers. Given these challenges, there is a growing need for adaptive RAG systems that dynamically adjust retrieval strategies based on query complexity Mallen et al. (2023).
@@
Adaptive-RAG Jeong et al. (2024) approach have attempted to balance accuracy and retrieval cost by inferring question complexity with a classifier to categorize queries into the most suitable retrieval strategies (e.g., no retrieval, single-step retrieval, or multi-step retrieval). it remains impractical due to its inability to dynamically adjust the accuracy-cost trade-off. Specifically, it lacks user-driven flexibility, preventing fine-grained control over retrieval strategies to suit diverse application needs. In real-world scenarios, retrieval preferences vary depending on the task domain, computational constraints, and user priorities—some applications may prioritize minimal cost, while others require higher accuracy despite increased retrieval overhead.
@@
To overcome these limitations, we propose Flare-Aug (FLexible, Adaptive REtrieval-Augmented Generation), a user-controllable RAG framework that enables fine-grained adjustment of the cost-accuracy trade-off. Flare-Agu introduces two classifiers: one trained to prioritize accuracy and another trained to prioritize cost. To seamlessly balance these two objectives, Flare-Aug incorporates an interpretable and easy-to-tune parameter $\alpha$, which controls the weighting between the two classifiers. This simple yet effective approach allows users to control their own accuracy and cost trade-off that best align with their specific requirements and user applications.
2 Relative Work
@@
Several works focus on leveraging LLM internal states to adapt retrieval dynamically. Baek et al. (2024) utilize hidden state representations from intermediate layers of language models to determine whether additional retrieval is necessary for a given query. Similarly, Yao et al. (2024) extract self-aware uncertainty from LLMs’ internal states to decide when retrieval should be performed. Jiang et al. (2023) propose an iterative retrieval approach, where the model predicts the upcoming sentence and uses it as a retrieval query if the sentence contains low-confidence tokens. However, these methods require access to the LLM’s internal states, making them impractical for many real-world applications, especially given the increasing reliance on proprietary, closed-source LLMs.
Another line of work trains external classifiers to predict retrieval necessity. Wang et al. (2023) train a classifier for LLM self-knowledge, allowing models to switch adaptively between retrieval and direct response generation. Jeong et al. (2024) introduce an adaptive retrieval framework that dynamically selects among no retrieval, single-step retrieval, or multi-step retrieval based on query complexity. Tang et al. (2024) propose a multi-arm bandit-based approach, where the model explores different retrieval strategies and optimizes retrieval choices based on feedback. Wang et al. (2024) develop an adaptive retrieval framework for conversational settings, determining whether retrieval is necessary based on dialogue context. Mallen et al. (2023) take an entity-centric approach, retrieving only when the entity popularity in a query falls below a certain threshold.
Despite these advancements, none of these approaches enable direct user control over retrieval strategies.
3 Problem Setting
3.1 Diverse User Query Complexities and Retrieval Strategies
@@ 3.2 Adaptive Retrieval
While previous works, such as Adaptive-RAG Jeong et al. (2024), have attempted to balance efficiency and cost by training classifiers using data collected from model predictions and the inductive dataset biases, they are adaptive only to static query complexities but not to user preferences, making them insufficient for handling the diverse needs of real-world applications. A truly adaptive retrieval system should incorporate both query complexity and user-level controllability. In addition to diverse query complexities, different users and application settings may prioritize accuracy over cost or vice versa, depending on their specific requirements. For example, a medical researcher using an LLM to summarize clinical trial results may prefer higher accuracy and completeness, even at the cost of increased retrieval latency and computational expenses. In contrast, a customer service chatbot handling frequent, simple user queries (e.g., ”What is the return policy?”) might prioritize fast, cost-effective responses, as occasional inaccuracies pose only negligible risk. Similarly, a stock market analyst requiring precise, up-to-date financial information might opt for a retrieval-intensive approach, despite longer response times, whereas a real-time virtual assistant answering everyday factual queries should minimize retrieval overhead, emphasizing speed over exhaustive accuracy. Given these diverse application domains and varying user demands, a one-size-fits-all adaptive retrieval approach is still suboptimal. Instead, a truly adaptive RAG system should be both efficient and customizable, enabling dynamically optimize the trade-off between accuracy and cost while allowing flexible user control.
@@ 4 Flare-Aug: Flexible Adaptive Retrieval Augmented Generation
In this section, we present the main component of our framework, which contains two classifier, Cost-Optimized Classifier, which is dynamic and LLM-dependent, and Reliability-Optimized Classifier, which is static and dataset-dependent. The overview of the Flare-Aug framework is illustrated in Figure 2.
@@ 4.1 Cost-Optimized Classifier
The Cost-Optimized Classifier selects the cheapest retrieval strategy that still ensures a correct answer. This classifier is LLM-dependent, since the training data is collected specifically from the LLM used for retrieval, and the prediction is sensitive and tailored to the LLM. The query in the training data is created by using the LLM to answer with all the three strategies, for each query, among the correctly answered strategies, using the strategy with minimal cost as the gold label. Thus, the labeling process itself is biased towards maximally eliminating unnecessary retrieval. The trained classifier, as a result, minimizes retrieval cost, latency, and computational complexity while maintaining an acceptable level of accuracy. However, a key concern of this classifier is that LLM may arrive at the correct answer by chance or by exploiting shortcuts in reasoning, leading to unverified accuracy. While this classifier is highly beneficial for cost-sensitive and efficiency-driven applications, it is unsuitable for high-risk domains where retrieval must be systematic and reliably grounded in external knowledge.
@@
The labeling process for the cost-optimized classifier is straightforward: we aim to select the cheapest retrieval strategy that still ensures a correct answer. Formally, let $S=\{s_{no},s_{sg},s_{multi}\}$ denote the set of available retrieval strategies: no retrieval ($s_{no}$), single-step retrieval ($s_{sg}$) and multi-step retrieval $(s_{multi})$. For each query $q$ in the training data, we evaluate whether it can be answered correctly under each retrieval strategy by assessing the LLM’s correctness with these strategies. Denote $LLM(q,s)$ as the correctness function, where:
$LLM(q,s)=\begin{cases}1,&\text{if the LLM answers correctly with retrieval strategy }s,\\ 0,&\text{otherwise}.\end{cases}$
@@ 4.2 Reliability-Optimized Classifier
The Reliability-Optimized Classifier assigns retrieval strategies based on intrinsic bias of the question answering, ensuring consistency, stability and reliability across different LLMs. This classifier are trained with only binary label: single-step retrieval and multi-step retrieval, with data from both single-hop QA dataset and multi-hop QA dataset. The accurate prediction of these dataset relies on the assumption that, single hop QA dataset and multi-hop QA dataset are intrinsically biases towards single-step retrieval and multi-step retrieval, respectively, and thus, may provide signals indicating which strategy to use during retrieval. Since there are only two labels, this classifier guarantees to have at least one retrieval, it may induce unnecessary computational overhead when using more powerful LLMs or handling straightforward queries that do not require retrieval. Despite this, since it is not LLM specific, it is more stable and reliable, encouraging higher accuracy and more reliable response. This classifier is effective in applications where accuracy is paramount, albeit at a higher cost.
@@
The Reliability-Optimized Classifier is trained using binary labels—single-step retrieval and multi-step retrieval—derived from question-answering datasets. Unlike the Cost-Optimized Classifier, which tailors retrieval strategy to a specific LLM, this classifier relies on dataset-level biases to infer retrieval needs, making it more stable across different LLMs. For each query $q$, we assign a label $s^{*}(q)$ based on the dataset it belongs to:
@@ 5.1 Experimental Setups
To assess the effectiveness of our approach in balancing accuracy, cost, and user control flexibility, we compare it against both adaptive and static retrieval strategies. Specifically, we evaluate our method against the adaptive retrieval approach Adaptive-RAG Jeong et al. (2024), as well as three static retrieval strategies: always using No Retrieval; always using Single-Step Retrieval and always using Multi-Step Retrieval (Trivedi et al., 2022a).
@@
We evaluate the trade-off between accuracy and retrieval cost using two key metrics. Accuracy measures whether the predicted answer contains the ground-truth answer. Retrieval cost is quantified as the number of retrieval steps required to generate an answer, where No Retrieval has a cost of 0, Single-Step Retrieval has a cost of 1, and Multi-Step Retrieval incurs a variable cost depending on the number of retrieval steps per query. By jointly considering these two metrics, we assess how different retrieval strategies balancing the trade-off between the response quality and computational efficiency.
@@
Following Jeong et al. (2024) and Trivedi et al. (2022a), we use BM25, a term-based sparse retrieval model, as the retriever. For queries from multi-step datasets, we utilize the pre-processed corpus from Trivedi et al. (2022a) as the external document corpus, while for queries from single-hop datasets, we use Wikipedia, preprocessed by Karpukhin et al. (2020), as the external document corpus. For answer generation, we employ two sizes of Flan-T5 models (Chung et al., 2024) (Flan-T5 XL and Flan-T5 XXL) and two sizes of GPT-4o models Hurst et al. (2024) (GPT-4o and GPT-4o-mini). For the classifier, we train a T5-Large model Raffel et al. (2020) using 500 samples from each dataset, resulting in a total of 3,000 queries for the labeling process. The classifier is trained with a learning rate of 3e-5 and a batch size of 64 on a single 80GB A100 GPU. We train the model for 20 epochs, as the loss has converged stably without indications of overfitting. We demonstrate the robustness of the classifier in Appendix A, with experiments on different training epochs and different sizes of classifiers.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has emerged as a powerful approach to mitigate large language model (LLM) hallucinations by incorporating external knowledge retrieval. However, existing RAG frameworks often apply retrieval indiscriminately, leading to inefficiencies—over-retrieving when unnecessary or failing to retrieve iteratively when required for complex reasoning. Although recent retrieval strategies can adaptively navigate among alternative retrieval strategies, they make their selection based solely on query complexity and incorporate no mechanism for prioritizing speed over accuracy or vice versa. This lack of user-defined control makes their use infeasible for diverse user application needs. In this paper, we introduce a novel user-controllable RAG framework that enables dynamic adjustment of the accuracy-cost trade-off. Our approach leverages two classifiers: one trained to prioritize accuracy and another to prioritize retrieval efficiency. Via an interpretable control parameter $\alpha$, users can seamlessly navigate between minimal-cost retrieval and high-accuracy retrieval depending on their specific requirements. We empirically demonstrate that our approach effectively balances accuracy, retrieval cost, and user controllability  [^1] Code is available at https://github.com/JinyanSu1/Flare-Aug., making it a practical and adaptable solution for real-world applications.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Khandelwal et al., 2019; Izacard et al., 2023) has emerged as a promising approach to address large language models’ (LLM) hallucinations, outputs that appear accurate but are actually factually incorrect. Hallucinations occur more often when the facts necessary to answer a query are outside the scope of the model’s training either because the facts are too recent, too obscure or answered by proprietary data (Trivedi et al., 2022a; Yao et al., 2022). By integrating retrieval modules, RAG enables LLMs to access and incorporate relevant external information, helping them stay updated with evolving world knowledge and reducing hallucination.
@@
Early work on retrieval-augmented LLMs focused primarily on single-hop queries (Lazaridou et al., 2022; Ram et al., 2023), where a single retrieval step retrieves relevant information based solely on the query. However, many complex queries require multi-step reasoning to arrive at the correct answer. For instance, a question such as, ”When did the people who captured Malakoff come to the region where Philipsburg is located?”, cannot be answered with a single retrieval step. Instead, such queries require an iterative retrieval process, in which the model retrieves partial knowledge, reasons over it, and conducts additional retrievals based on intermediate conclusions (Trivedi et al., 2022a; Press et al., 2022; Jeong et al., 2024). While multi-step retrieval enables deeper reasoning, it also incurs significant computational overhead. Conversely, for simple queries, LLMs may already encode sufficient knowledge to generate an accurate response without any retrieval whatsoever. However, many existing RAG frameworks apply retrieval indiscriminately, either over-relying on external knowledge when parametric memory suffices or failing to retrieve iteratively when deeper reasoning is required. This lack of adaptability results in inefficiencies: unnecessary retrieval increases latency and computational cost, while insufficient retrieval leads to incomplete or incorrect answers. Given these challenges, there is a growing need for adaptive RAG systems that dynamically adjust retrieval strategies based on query complexity (Mallen et al., 2023).
@@
The Adaptive-RAG (Jeong et al., 2024) approach attempted to balance accuracy and retrieval cost by inferring question complexity via a classifier that categorizes a query according to its most suitable retrieval strategy (e.g., no retrieval, single-step retrieval, or multi-step retrieval). But the method remains impractical due to its inability to dynamically adjust the accuracy-cost trade-off. Specifically, it lacks user-driven flexibility, preventing fine-grained control over retrieval strategies in order to support diverse application needs. In real-world scenarios, retrieval preferences vary depending on the task domain, computational constraints, and user priorities: some applications may prioritize (low) cost, while others require higher accuracy despite increased retrieval overhead.
@@
To overcome these limitations, we propose Flare-Aug (FLexible, Adaptive REtrieval-Augmented Generation), a user-controllable RAG framework that enables fine-grained adjustment of the cost-accuracy trade-off. Flare-Agu introduces two classifiers: one trained to prioritize accuracy and another trained to prioritize cost. To seamlessly balance these two objectives, Flare-Aug incorporates an interpretable and easy-to-tune parameter $\alpha$, which controls the weighting between the two classifiers. This simple yet effective approach allows users to control their own accuracy and cost trade-off to best align with their specific requirements and user applications.
2 Related Work
@@
Several works focus on leveraging LLM internal states to adapt retrieval dynamically. Baek et al. (2024) utilize hidden state representations from intermediate layers of language models to determine whether additional retrieval is necessary for a given query. Similarly, Yao et al. (2024) extract self-aware uncertainty from LLMs’ internal states to decide when retrieval should be performed. And Jiang et al. (2023) propose an iterative retrieval approach, where the model predicts the upcoming sentence and uses it as a retrieval query if the sentence contains low-confidence tokens. However, these methods require access to the LLM’s internal states, making them impractical for many real-world applications, especially given the increasing reliance on proprietary, closed-source LLMs.
Another line of work trains external classifiers to predict retrieval necessity. Wang et al. (2023) train a classifier for LLM self-knowledge, allowing models to switch adaptively between retrieval and direct response generation. Jeong et al. (2024) introduce an adaptive retrieval framework that dynamically selects among no retrieval, single-step retrieval, or multi-step retrieval based on query complexity. Tang et al. (2024) propose a multi-arm bandit-based approach, where the model explores different retrieval strategies and optimizes retrieval choices based on feedback. Wang et al. (2024b) develop an adaptive retrieval framework for conversational settings, determining whether retrieval is necessary based on dialogue context. Mallen et al. (2023) take an entity-centric approach, retrieving only when the entity popularity in a query falls below a certain threshold.
Despite these advancements, none of these approaches enables direct user control over retrieval strategies.
3 Problem Setting: Diverse User Query Complexities and Retrieval Strategies
@@ 3.1 Adaptive Retrieval
While previous work, such as Adaptive-RAG (Jeong et al., 2024), has attempted to balance efficiency and cost by training classifiers using data collected from model predictions and the inductive biases in the datasets themselves, they are adaptive only to static query complexities but not to user preferences, making them insufficient for handling the diverse needs of real-world applications. A truly adaptive retrieval system should incorporate both query complexity and user-level controllability. In addition to diverse query complexities, different users and application settings may prioritize accuracy over cost or vice versa, depending on their specific requirements. For example, a medical researcher using an LLM to summarize clinical trial results may prefer higher accuracy and completeness, even at the cost of increased retrieval latency and computational expenses. In contrast, a customer service chatbot handling frequent, simple user queries (e.g., ”What is the return policy?”) might prioritize fast, cost-effective responses, as occasional inaccuracies pose only negligible risk. Similarly, a stock market analyst requiring precise, up-to-date financial information might opt for a retrieval-intensive approach, despite longer response times, whereas a real-time virtual assistant answering everyday factual queries should minimize retrieval overhead, emphasizing speed over exhaustive accuracy. Given these diverse application domains and varying user demands, a one-size-fits-all adaptive retrieval approach is still suboptimal. Instead, a truly adaptive RAG system should be both efficient and customizable, enabling dynamically optimized trade-offs between accuracy and cost while allowing flexible user control.
@@ 4 Flare-Aug: Flexible Adaptive Retrieval Augmented Generation
In this section, we present the main components of our framework, a Cost-Optimized Classifier, which is dynamic and LLM-dependent, and a Reliability-Optimized Classifier, which is static and dataset-dependent. An overview of the Flare-Aug framework is illustrated in Figure 2 and described in the sections below.
@@ 4.1 Cost-Optimized Classifier
The goal of the Cost-Optimized Classifier is to select for a given query the cheapest retrieval strategy that still ensures a correct answer. This classifier is LLM-dependent, since the training data is collected specifically from the LLM used for retrieval, and the prediction is sensitive and tailored to the LLM. The $<$query, retrieval strategy$>$ pairs that comprise the training data are created by using the LLM to answer with all the three strategies–no retrieval, single-step retrieval and multi-step retrieval: the gold label for each query is the minimal cost strategy that produces a correct answer. Thus, the labeling process itself is biased towards maximally eliminating unnecessary retrieval. The trained classifier, as a result, minimizes retrieval cost, latency, and computational complexity while maintaining an acceptable level of accuracy. (See the experiments in later section when $\alpha=0$). However, a key concern for this classifier is that the LLM may arrive at the correct answer by chance or by exploiting shortcuts in reasoning, leading to unverifiable answers. While this classifier is highly beneficial for cost-sensitive and efficiency-driven applications, it is unsuitable for high-risk domains where retrieval must be systematic and reliably grounded in external knowledge.
@@
The labeling process for the cost-optimized classifier is straightforward: we aim to select the cheapest retrieval strategy that still ensures a correct answer. Formally, let $S=\{s_{no},s_{sg},s_{multi}\}$ denote the set of available retrieval strategies: no retrieval ($s_{no}$), single-step retrieval ($s_{sg}$) and multi-step retrieval $(s_{multi})$. For each query $q$ in the training data, we evaluate whether the LLM answers it correctly under each retrieval strategy. Denote $LLM(q,s)$ as the correctness function, where:
$LLM(q,s)=\begin{cases}1,&\text{if the LLM answers $q$ correctly with retrieval strategy }s,\\ 0,&\text{otherwise}.\end{cases}$
@@ 4.2 Reliability-Optimized Classifier
The goal for the Reliability-Optimized Classifier is to assign a retrieval strategy to a query based on intrinsic difficulty of the query itself, ensuring consistency, stability and reliability across different LLMs. This classifier is trained on $<$query, retrieval strategy$>$ pairs with binary labels: single-step retrieval and multi-step retrieval, using data from a single-hop QA dataset and multi-hop QA dataset, respectively. Here we rely on the assumption that the single hop QA dataset is in fact intrinsically biased towards questions that require single-step retrieval, and the multi-hop QA dataset towards questions that require multi-step retrieval. Since there are only two labels, this classifier guarantees least one retrieval, and may induce unnecessary computational overhead when using more powerful LLMs or handling straightforward queries that do not require retrieval. Despite this, since it is not LLM specific, it is more stable and reliable, encouraging higher accuracy and more reliable response. (See experiments in later section when $\alpha=1$). This classifier is effective in applications where accuracy is paramount, albeit at a higher cost.
@@
The Reliability-Optimized Classifier is trained using binary labels—single-step retrieval and multi-step retrieval—derived from existing question-answering datasets. Unlike the Cost-Optimized Classifier, which tailors the retrieval strategy to a specific LLM, this classifier relies on dataset-level biases to infer retrieval needs, making it more stable across different LLMs. For each query $q$, we assign a label $s^{*}(q)$ based on the dataset it belongs to:
@@ 5.1 Experimental Setup
To assess the effectiveness of our approach in balancing accuracy, cost, and user control flexibility, we compare it against both adaptive and static retrieval strategies. Specifically, we evaluate our method against the adaptive retrieval approach Adaptive-RAG (Jeong et al., 2024), as well as three static retrieval strategies: always use No Retrieval; always use Single-Step Retrieval and always use Multi-Step Retrieval (Trivedi et al., 2022a).
@@
We evaluate the trade-off between accuracy and retrieval cost using two metrics. Accuracy measures whether the predicted answer contains the ground-truth answer. Retrieval cost is quantified as the number of retrieval steps required to generate an answer, where No Retrieval has a cost of 0, Single-Step Retrieval has a cost of 1, and Multi-Step Retrieval incurs a variable cost depending on the number of retrieval steps per query. By jointly considering these two metrics, we assess how different retrieval strategies balance the trade-off between response quality and computational efficiency. Note that we use retrieval-side cost as a proxy for cost due to its Interpretability and simplicity: Retrieval step count is a transparent and easily measurable proxy that correlates with latency, token consumption, and compute in RAG systems. Unlike in math reasoning tasks—where reasoning token length is often used as a proxy for cost—the context length in RAG is not a reliable cost metric. This is because context size in RAG is largely dictated by the retrieval configuration (e.g., top-k) rather than the intrinsic complexity of the query. For example, retrieving 10 passages versus 2 directly affects context length without reflecting deeper reasoning complexity. In Appendix A.1, we further justified the choice of these two evaluation metrics.
@@
Following Jeong et al. (2024) and Trivedi et al. (2022a), we use BM25, a term-based sparse retrieval model, as the retriever [^2] While BM25 is a sparse retriever, it remains a strong and widely-used baseline in the adaptive RAG approaches such as Jeong et al. (2024); Tang et al. (2024); Jiang et al. (2023); Su et al. (b); Wang et al. (2024a; 2023); Trivedi et al. (2022a); Mallen et al. (2023). Studies (e.g., Ram et al. (2023)) show that BM25 continues to outperform many dense retrievers.. For queries from multi-step datasets, we utilize the pre-processed corpus from Trivedi et al. (2022a) as the external document corpus, while for queries from single-hop datasets, we use Wikipedia, preprocessed by Karpukhin et al. (2020), as the external document corpus. For answer generation  [^3] We used the same prompting format as in prior work Trivedi et al. (2022a); Jeong et al. (2024), the exact prompts we used are publicly available in the codebase associated with Trivedi et al. (2022a), which can be accessed at: https://github.com/StonyBrookNLP/ircot/tree/main, we employ two sizes of Flan-T5 models (Chung et al., 2024) (Flan-T5 XL and Flan-T5 XXL) and two sizes of GPT-4o models (Hurst et al., 2024) (GPT-4o and GPT-4o-mini). For the classifier, we train a T5-Large model (Raffel et al., 2020) using 500 samples from each dataset, resulting in a total of 3,000 queries for the labeling process. The classifier is trained with a learning rate of 3e-5 and a batch size of 64 on a single 80GB A100 GPU. We train the model for 20 epochs, as the loss has converged stably without indications of overfitting  [^4] Both classifiers are lightweight and fast to train. For instance, training the Cost-Optimized Classifier on GPT-4o-mini–specific data (1,377 samples,  450 steps) takes only 223 seconds with a batch size of 64. The Reliability-Optimized Classifier (trained on 2,400 samples,  750 steps) takes 417 seconds. In total, training both classifiers requires approximately 640 seconds. We refer readers to discussions in Appendix A.2.. We demonstrate the robustness of the classifier in Appendix B, with experiments on different training epochs and different sizes of classifiers.